OpenStudy (anonymous):

A first number plus twice a second number is 24. Twice the first number plus the second number is 21. What are the numbers?

OpenStudy (haleyelizabeth2017):

x+2y=24 is going to be the equation for the first one....

OpenStudy (haleyelizabeth2017):

Can you try to find the second one?

OpenStudy (anonymous):

y+2x=21

OpenStudy (haleyelizabeth2017):

=24*

OpenStudy (haleyelizabeth2017):

but I would word it to where the x term comes first so it is easier :)

OpenStudy (anonymous):

oh I typed my first one in wrong, it is actually 21

OpenStudy (haleyelizabeth2017):

oh okay :)

sammixboo (sammixboo):

Correct, and \(\huge \rlap{\bbox[2.9pt, ,border:6.8px solid #2aacd7 ]{\Huge\color{blueviolet}{\rm \fbox{Welcome}}}}\hspace{127pt}\huge \rlap{\bbox[3.3pt, ,border:7px solid #2aacd7 ]{\Huge\color{blueviolet}{\rm \fbox{to}}}}\) \(\huge \color{blueviolet}{\heartsuit^{\LARGE \heartsuit^{\Large \heartsuit}}}\) \(\huge \color{blueviolet}{\heartsuit^{\LARGE \heartsuit^{\Large \heartsuit}}}\) \(\huge \huge \rlap{\bbox[5pt, ,border:8px solid #2aacd7 ]{\Huge\color{#84c73c}{\rm \fbox{OpenStudy}}}}\) If you need help using this site you can just ask me or any user. Have a good day! If you need a second opinion also, I am here

OpenStudy (haleyelizabeth2017):

Are they both 21 or is one of them 24 still?

OpenStudy (anonymous):

One of them is 24 I edited the original question and fixed it.

OpenStudy (haleyelizabeth2017):

okay great! So now that we have both: x+2y=24 2x+y=21 We can now substitute to find the values of x and y :) I can show you the steps if you want?

OpenStudy (anonymous):

Ok, I tried a lot of the different combinations of numbers that fit for the first one, but none of them worked for the second one, they were all way too big.

OpenStudy (haleyelizabeth2017):

Substitution is not where you replace numbers into x and y, but rather one equation into the other. So first, we need to isolate x for the first one.

OpenStudy (anonymous):

Ok so x=24-2y?

OpenStudy (haleyelizabeth2017):

Correct! So now that we have that, we can then plug it into the x of the second equation like so: 2(-2y+24)+y=21 Then you need to distribute the 2 to the -2y and 24.

OpenStudy (haleyelizabeth2017):

Do you understand what I did there?

OpenStudy (anonymous):

-4y+48+y=21

OpenStudy (haleyelizabeth2017):

Correct! Then you need to combine like terms.

OpenStudy (anonymous):

which can be simplified to -3y=-27

OpenStudy (haleyelizabeth2017):

Correct! Then divide both sides by -3

OpenStudy (anonymous):

so y=9

OpenStudy (anonymous):

so put that into the next equation to solve for x

OpenStudy (haleyelizabeth2017):

Yes. Then you can replace y=9 into either one of the equations (It doesn't matter which) and solve

OpenStudy (haleyelizabeth2017):

It would be easier to put it into the x=-2y+24 since x is already isolated

OpenStudy (anonymous):

x+18=24, so x equals 6!

OpenStudy (anonymous):

Thank you so much for the help! :) :)

OpenStudy (haleyelizabeth2017):

Yes! Great job! So: y=9 x=6 And that's all! :D

OpenStudy (haleyelizabeth2017):

No problem!

OpenStudy (haleyelizabeth2017):

Would you mind hitting the 'close' button to close this question? Thank you!

OpenStudy (anonymous):

Ok

OpenStudy (haleyelizabeth2017):

Thank you!